OpenStudy (anonymous):

Find the value of the discriminant for the following equation: x^2 - 9x + 14 = 0

OpenStudy (anonymous):

lawdamercy \[9^2-4\times 1\times 14=81-56=25\]

OpenStudy (anonymous):

that be \[b^2-4ac\] for \[b=-9,a=1,c=14\]

OpenStudy (anonymous):

Thank you ghun